OpenStudy (anonymous):

find LCM...10b^4c^9, 2b^6c^8, 15b^7c^6

OpenStudy (anonymous):

ok im confused on how to write the problem out. The way I did it was \[2\times5.b.b.b.b.b.b.b.b.b.b.b.b.b\]

OpenStudy (anonymous):

the way i did it was what two numbers equal 10 (2x5) then i thought we write out how many b's or c's there are

hero (hero):

Forget what I wrote previously.

hero (hero):

Instead, think of trying to find the Least Common Denominator

OpenStudy (anonymous):

i did that with each number

hero (hero):

\[\frac{1}{10b^4c^9}+\frac{1}{2b^6c^8}+\frac{1}{15b^7c^6}\] Imagine you had to add those up. How would you find the LCD?

OpenStudy (anonymous):

would that be 2?

hero (hero):

30b^7c^9 is the correct answer here.

OpenStudy (anonymous):

ok i have no clue how you got that lol

OpenStudy (anonymous):

so the way i was working it, isnt a good way?